Briannacook19:

Angel Graham takes out a $150,000 mortgage. This is a 30-year at $725 per month. What is the total amount of interest that Angel will pay on this mortgage? Total Interest = $_____

Ultrilliam:

First we need to convert 30 years to months, 12*30=x Next we need to multiply that number by the monthly payment, so 725*x=y Finally to get the total amount of interest we need to subtract the original number (150,000) from the result of the before problem, so y-150,000=z That said, doing the math, whats the result?